# Efficient and Comprehensive Feature Extraction in Large Vision-Language Model for Clinical Pathology Analysis

### v2

@@ 4.1 Mixed Task-Guided Feature Enhancement
Currently, pathology LVLMs [40, 36, 31, 13] commonly use the vision encoder from Contrastive Language-Image Pre-Training (CLIP) [35] to convert images into embedding tokens. Although this vision encoder enables alignment between image features and the text space, facilitating the LLM’s understanding of image content, it primarily relies on pre-training data consisting of image-caption pairs. This leads the vision encoder to excel at extracting global features of images but limits its ability to perceive local details and spatial structures. However, in pathological diagnosis, accurately identifying foundational pathological concepts and their spatial relationships within pathology WSIs is essential. For instance, in diagnosing microvascular invasion [54], a pathologist needs to first locate the cancerous region and then inspect surrounding vessels for the presence of cancer cell nuclei, requiring the model to have a nuanced understanding of pathological concepts such as cancerous tissue, blood vessels, normal cell nuclei, and cancer cell nuclei, along with their spatial relationships. Current pathology LVLMs still exhibit limitations in this regard. To address this issue, we focus on both training data and model architecture to enhance the model’s capability in extracting and understanding detailed features.
@@
Table 1: Performance comparison on patch-level pathology diagnostic tasks.
To further improve the model’s ability to recognize multidimensional pathological features, we also constructed a cross-scale instruction task set that includes organ recognition, cancer subtype identification and grading, tissue type recognition, microsatellite instability detection, and tumor-infiltrating lymphocyte recognition. These self-constructed datasets, combined with publicly available pathology visual question-answering datasets, have formed a training dataset containing 20 types of organs and approximately 490,000 training samples, significantly enhancing the model’s ability to extract pathological features at multiple scales and granularities. For details on dataset sources and construction methods, please refer to the supplementary material.
In terms of model architecture, we implemented three primary improvements (as shown in Figure 3). First, we added an extra ViT named UNI [11]. This ViT, pretrained on large-scale pathology images using the DINOv2 [33] framework, provides critical fine-grained pathological visual features for detection and segmentation tasks. Second, we adopted a multi-scale feature fusion strategy, enabling the model to handle higher-resolution input images without retraining the vision encoder. Specifically, we set a series of increasing input resolutions based on the original resolution supported by the vision encoder, with other resolutions as integer multiples of this base. For the original resolution images, features are directly extracted via the vision encoder. For higher-resolution images, they are divided into image tiles of the original size, each processed separately for feature extraction, and then reassembled based on spatial position. Finally, the features are averaged to match the original feature map dimensions and concatenated along the channel dimension. This approach allows for richer detail extraction without increasing the number of input image tokens $e_{v}$ to the LLM $\mathcal{M}$.
@@
Table 2: Performance comparison on slide-level pathology diagnostic tasks.
Metric Cancer Region Tissue Structure Cell Nucleus
HCC ICC RCC GBM LUAD BC LD VD ND MoNuSeg NuCLS PanNuke
F1-Score 92.13 100.00 90.91 40.85 90.91 41.38 82.70 89.52 72.50 83.98 34.30 45.30
IoU 91.97 91.85 89.59 73.85 87.96 72.75 87.34 83.00 79.75 79.65 44.98 59.99
Table 3: Performance of OmniPath in pathology referring detection tasks.
4.2 Prompt-Guided Detail Feature Completion
When using pathology LVLMs for WSI diagnostic analysis, only the thumbnail of the WSI can be used as input, resulting in substantial information loss that affects diagnostic accuracy. Our proposed PGDFC strategy dynamically completes missing information based on specific task requirements while maintaining inference efficiency, as shown in Figure 3. Specifically, we first remove elements corresponding to image background regions from $\Psi_{(N+M),e_{v}}$ and select the top-$S$ elements with the highest values from the remaining elements, denoting their indices as $\mathbf{I}=\{i_{1},i_{2},\dots,i_{S}\}$. Then, using the index set $\mathbf{I}$, we locate the corresponding tile regions in the original WSI and extract high-resolution patches from these regions, denoted as $\hat{x}_{v}=\{\hat{x}_{v}^{1},\hat{x}_{v}^{2},\dots,\hat{x}_{v}^{S}\}$.
We use $\mathcal{V}$ to extract features for each image in $\hat{x}_{v}$, obtaining a set of feature sequences $\hat{e}_{v}=\{\hat{e}_{v}^{1},\hat{e}_{v}^{2},\dots,\hat{e}_{v}^{S}\}$, where $\hat{e}_{v}^{s}=\mathcal{V}(\hat{x}_{v}^{s})$. Simultaneously, we encode the positional information of each patch using textual descriptions, denoted as $\hat{x}_{t}=\{\hat{x}_{t}^{1},\hat{x}_{t}^{2},\dots,\hat{x}_{t}^{S}\}$, and obtain the corresponding text embeddings $\hat{e}_{t}=\{\hat{e}_{t}^{1},\hat{e}_{t}^{2},\dots,\hat{e}_{t}^{S}\}$ through $\mathcal{T}$, where $\hat{e}_{t}^{s}=\mathcal{T}(\hat{x}_{t}^{s})$. By feeding $\hat{e}_{v}$ and $\hat{e}_{t}$ along with the original inputs $e_{v}$ and $e_{t}$ into $\mathcal{M}$, we can obtain the final diagnostic result $y=\mathcal{M}(e_{v},e_{t},\hat{e}_{v},\hat{e}_{t})$. To mitigate the impact of the large number of $\hat{e}_{v}$ tokens on $\mathcal{M}$’s inference efficiency, we use the average pooling on each element of $\hat{e}_{v}$ to reduce the token count of it.
5 Experiments
In this section, we first provide the implementation details of OmniPT, followed by comparative results across multiple tasks. Finally, we conduct ablation studies on the key components of OmniPT.
5.1 Implementation Details
Based on the pretrained LLaVA-1.5-13B [30], we constructed OmniPT by replacing its CLIP ViT-L 336px [35] visual encoder with SigLIP ViT-SO 384px [53] and integrating UNI [11] as an auxiliary vision encoder. The projector utilizes a two-layer MLP with GELU activation. The mask encoder is implemented with ResNet-50 [19], while the mask decoder follows SAM’s [24] decoder architecture but directly uses the LVLM’s vision encoder to replace SAM’s original image encoder for feature extraction. During training, all modules of OmniPT participate in end-to-end training with no parameter freezing. We created a multitask dataset encompassing 20 organs and approximately 490,000 samples for model training (see supplementary material for detailed data sources and construction methods). Unlike the two-stage training strategy commonly adopted by existing LVLMs, OmniPT requires only a single-stage fine-tuning: trained for 2 epochs on 8 NVIDIA A800 GPUs using the AdamW optimizer, with a learning rate of 2e-5 and a global batch size of 128. We set $S=8$, and compare OmniPT with LLaVA-1.5 [30], LLaVA-Med [27], Quilt-LLaVA [36], and PA-LLaVA [13].
5.2 Comparison on Pathology Diagnostic Tasks
To evaluate the pathology diagnostic performance of various LVLMs, we conducted a series of clinically relevant pathology diagnostic experiments, divided by diagnostic granularity into patch-level and slide-level categories. The patch-level experiments included subtyping and grading tasks for a range of common cancers, such as hepatocellular carcinoma subtyping (HCC-S) and grading (HCC-G), intrahepatic cholangiocarcinoma subtyping (ICC-S) and grading (ICC-G), renal cell carcinoma subtyping (RCC-S), lung cancer subtyping (LUNG-S) and grading (LUNG-G), gastric adenocarcinoma Lauren subtyping (STAD-L) and grading (STAD-G). In addition, other tasks related to pathology concept recognition or diagnosis included: microvascular invasion identification (MVI), neural invasion identification (NI), pan-cancer identification across 32 types (PanCancer), organ classification (OC), tissue classification (TC), tumor-infiltrating lymphocyte identification (TIL), microsatellite instability detection in colorectal cancer (MSI), and seven-class gastric lesion recognition (GLR-7). The accuracy of each model on these tasks is detailed in Table 1.
Slide-level experiments included not only the same subtyping and grading tasks as the patch-level but also additional tasks, such as lymph node metastasis diagnosis (LNM), HCC prognosis prediction (HCC-P), and colorectal cancer prognosis prediction (CRC-P), using WSI thumbnails as image input. The accuracy of each model on slide-level tasks is presented in Table 2.
It can be observed that OmniPT achieves the best performance across all patch-level and slide-level pathological diagnosis tasks. For most cancer subtype classification and grading tasks, OmniPT achieves accuracy rates exceeding 70%, and in many cases, surpassing 90%. In contrast, the accuracy rates of other models are generally below 70%. This demonstrates that OmniPT is more suitable for clinical applications to assist pathologists in diagnosis. Furthermore, OmniPT also exhibits strong recognition capabilities for features such as microvascular invasion, neural invasion, and tumor-infiltrating lymphocytes. This significantly reduces the extensive effort required by pathologists to meticulously examine detailed pathological lesions during slide review.
@@ 5.3 Comparison on Zero-Shot Classification Tasks
To evaluate the clinical generalization capability of OmniPT, we employed a zero-shot classification paradigm, testing on several widely-used academic pathology datasets that were not included in the training set. The evaluation covered two levels: patch-level tasks using the CCRCC [6], MHIST [47], and NCT-CRC [32] datasets, and slide-level tasks using the PANDA [7], DHMC [48], and CAMELYON17 [29] datasets. Using a closed-ended question-answering approach, the model was required to classify images into predefined categories specific to each dataset. The performance comparison of all models on these zero-shot test sets is presented in Table 4.
It is shown that OmniPT consistently outperforms other models across both patch-level and slide-level datasets in zero-shot classification tasks, highlighting its strong generalization ability in pathological image analysis. Notably, on the PANDA and CAMELYON17 slide-level datasets, OmniPT achieved the highest accuracy rates of 79.15% and 59.33%, respectively, which significantly surpasses the performance of other models. This superior performance in zero-shot classification indicates OmniPT’s robustness in handling diverse pathological image data and reinforces its potential for clinical applications where labeled training data may be limited.
@@
Table 4: Comparison on zero-shot classification tasks.
Metric Cancer Region Tissue Structure
@@
IoU 91.65 88.25 80.03 39.34 87.08 64.16 81.31 54.01 50.96
Dice 95.51 93.60 85.11 50.01 93.03 74.64 89.23 62.14 56.94
Table 5: OmniPath’s performance in referring segmentation.
@@ 5.4 Detection and Segmentation Performance
In pathological diagnosis, detection and segmentation tasks are as critical as classification tasks. For instance, assessing microvascular invasion requires counting cancer cell nuclei within blood vessels, which relies on accurate referring detection. Given that existing pathological LVLMs lack detection and segmentation capabilities, this part presents only OmniPT’s performance on these tasks. The detection tasks cover various cancer region identifications, including HCC, ICC, RCC, glioblastoma (GBM), lung adenocarcinoma (LUAD), and bladder cancer (BC). Additionally, it involves detecting tissue structures such as lymph nodes (LD), vessels (VD), and nerves (ND), as well as cell nuclei detection in datasets like MoNuSeg [25] (without categories), NuCLS [3], and PanNuke [18] (with categories). The segmentation tasks not only include the segmentation of cancer regions covered by the detection tasks but also nerve segmentation (NS), nerve invasion region segmentation (NIS), and lymph node metastasis segmentation (LNMS). Detection tasks are evaluated using F1-score and IoU, while segmentation tasks are assessed using the IoU and Dice coefficient. Table 3 and Table 5 show the detection and segmentation performance, respectively.
While OmniPT’s performance may not surpass specialized smaller models on certain tasks, it offers a unique advantage in inferential capabilities. As shown in Figure 1, OmniPT can accurately detect cancer cell nuclei within blood vessels according to instructions—an ability that current specialized small models find challenging to achieve.

### v3

@@ 4.1 Mixed Task-Guided Feature Enhancement
Currently, pathology LVLMs [pathasst, quiltllava, pathchat, pallava] commonly use the vision encoder from Contrastive Language-Image Pre-Training (CLIP) [clip] to convert images into embedding tokens. Although this vision encoder enables alignment between image features and the text space, facilitating the LLM’s understanding of image content, it primarily relies on pre-training data consisting of image-caption pairs. This leads the vision encoder to excel at extracting global features of images but limits its ability to perceive local details and spatial structures. However, in pathological diagnosis, accurately identifying foundational pathological concepts and their spatial relationships within pathology WSIs is essential. For instance, in diagnosing microvascular invasion [Zhang_2023_CVPR], a pathologist needs to first locate the cancerous region and then inspect surrounding vessels for the presence of cancer cell nuclei, requiring the model to have a nuanced understanding of pathological concepts such as cancerous tissue, blood vessels, normal cell nuclei, and cancer cell nuclei, along with their spatial relationships. Current pathology LVLMs still exhibit limitations in this regard. To address this issue, we focus on both training data and model architecture to enhance the model’s capability in extracting and understanding detailed features.
To further improve the model’s ability to recognize multidimensional pathological features, we also constructed a cross-scale instruction task set that includes organ recognition, cancer subtype identification and grading, tissue type recognition, microsatellite instability detection, and tumor-infiltrating lymphocyte recognition. These self-constructed datasets, combined with publicly available pathology visual question-answering datasets, have formed a training dataset containing 21 types of organs and approximately 490K training samples, significantly enhancing the model’s ability to extract pathological features at multiple scales and granularities. For details on dataset sources and construction methods, please refer to the Appendix F.
In terms of model architecture, we implemented three primary improvements (as shown in Fig. 3). First, we added an extra ViT named UNI [UNI]. This ViT, pretrained on large-scale pathology images using the DINOv2 [dinov2] framework, provides critical fine-grained pathological visual features for detection and segmentation tasks. Second, we adopted a multi-scale feature fusion strategy, enabling the model to handle higher-resolution input images without retraining the vision encoder. Specifically, we set a series of increasing input resolutions based on the original resolution supported by the vision encoder, with other resolutions as integer multiples of this base. For the original resolution images, features are directly extracted via the vision encoder. For higher-resolution images, they are divided into image tiles of the original size, each processed separately for feature extraction, and then reassembled based on spatial position. Finally, the features are averaged to match the original feature map dimensions and concatenated along the channel dimension. This approach allows for richer detail extraction without increasing the number of input image tokens $e_{v}$ to the large language model $\mathcal{M}$.
Figure 3: Overview of the proposed OmniPath. Left: the architecture of OmniPath with the MGFE and PGFC strategy. The MGFE model module improvements include a multi-scale feature fusion vision encoder and an additional mask decoder. The PGFC process, shown by the red dashed line, involves inputting a WSI thumbnail and its corresponding prompt into OmniPath. The top-$S$ patches with the highest attention values are selected, and their higher-resolution images are retrieved from the original WSI and added as supplementary input to OmniPath. Right: the detailed structure of the multi-scale feature fusion vision encoder.
4.2 Prompt-Guided Detail Feature Completion
When using pathology LVLMs for WSI diagnostic analysis, only the thumbnail of the WSI can be used as input, resulting in substantial information loss that affects diagnostic accuracy. Our proposed PGFC strategy dynamically completes missing information based on specific task requirements while maintaining inference efficiency, as shown in Fig. 3. Specifically, we first remove elements corresponding to image background regions from $\Psi_{(N+M),e_{v}}$ and select the top-$S$ elements with the highest values from the remaining elements, denoting their indices as $\mathbf{I}=\{i_{1},i_{2},\dots,i_{S}\}$. Then, using the index set $\mathbf{I}$, we locate the corresponding tile regions in the original WSI and extract high-resolution patches from these regions, denoted as $\hat{x}_{v}=\{\hat{x}_{v}^{1},\hat{x}_{v}^{2},\dots,\hat{x}_{v}^{S}\}$.
We use $\mathcal{V}$ to extract features for each image in $\hat{x}_{v}$, obtaining a set of feature sequences $\hat{e}_{v}=\{\hat{e}_{v}^{1},\hat{e}_{v}^{2},\dots,\hat{e}_{v}^{S}\}$, where $\hat{e}_{v}^{s}=\mathcal{V}(\hat{x}_{v}^{s})$. In parallel, we encode the positional information of each patch using textual descriptions, denoted as $\hat{x}_{t}=\{\hat{x}_{t}^{1},\hat{x}_{t}^{2},\dots,\hat{x}_{t}^{S}\}$, and obtain the corresponding text embeddings $\hat{e}_{t}=\{\hat{e}_{t}^{1},\hat{e}_{t}^{2},\dots,\hat{e}_{t}^{S}\}$ through $\mathcal{T}$, where $\hat{e}_{t}^{s}=\mathcal{T}(\hat{x}_{t}^{s})$. By feeding $\hat{e}_{v}$ and $\hat{e}_{t}$ along with the original inputs $e_{v}$ and $e_{t}$ into $\mathcal{M}$, we can obtain the final diagnostic result $y=\mathcal{M}(e_{v},e_{t},\hat{e}_{v},\hat{e}_{t})$. To mitigate the impact of the large number of $\hat{e}_{v}$ tokens on $\mathcal{M}$’s inference efficiency, we use the average pooling on each element of $\hat{e}_{v}$ to reduce the token count of it.
5 Experiments
In this section, we first provide the implementation details of OmniPath, followed by comparative results across multiple tasks. Finally, we conduct ablation studies on the key components of OmniPath.
5.1 Implementation Details
Based on the pretrained LLaVA-1.5 [llava], we constructed OmniPath by replacing its CLIP ViT-L 336px [clip] visual encoder with SigLIP ViT-SO 384px [siglip] and integrating UNI [UNI] as an auxiliary vision encoder. The projector utilizes a two-layer MLP with GELU activation. The mask encoder is implemented with ResNet-18 [resnet], while the mask decoder follows SAM’s [sam] decoder architecture (randomly initialized instead of using pre-trained weights) but directly uses the LVLM’s vision encoder described above to replace SAM’s original image encoder for feature extraction. During training, all modules of OmniPath participate in end-to-end training with no parameter freezing. We created a multitask dataset encompassing 21 organs and approximately 490K samples for model training (see Appendix F for detailed data sources and construction methods). Unlike the two-stage training strategy commonly adopted by existing LVLMs, OmniPath requires only a single-stage fine-tuning: trained for 2 epochs on 8 NVIDIA A100 GPUs using the AdamW optimizer, with a learning rate of 2e-5 and a global batch size of 128. We set $S=8$, and compare OmniPath with LLaVA-1.5 [llava], LLaVA-Med [llavamed], Quilt-LLaVA [quiltllava], and PA-LLaVA [pallava].
For a fair comparison, we fine-tuned Quilt-LLaVA using the same training dataset and hyperparameters. The test results are shown in the table below as “Quilt-LLaVA (FT)”. Since its model architecture is identical to that of LLaVA, with only the weights differing, this result partially reflects the fine-tuning performance of other LLaVA-based methods.
5.2 Comparison on Pathology Diagnostic Tasks
To evaluate the pathology diagnostic performance of various LVLMs, we conducted a series of clinically relevant pathology diagnostic experiments, divided by diagnostic granularity into patch-level and slide-level categories. The patch-level experiments included subtyping and grading tasks for a range of common cancers, such as hepatocellular carcinoma subtyping (HCC-S) and grading (HCC-G), intrahepatic cholangiocarcinoma subtyping (ICC-S) and grading (ICC-G), renal cell carcinoma subtyping (RCC-S), lung cancer subtyping (LUNG-S) and grading (LUNG-G), gastric adenocarcinoma Lauren subtyping (STAD-L) and grading (STAD-G). In addition, other tasks related to pathology concept recognition or diagnosis included: microvascular invasion identification (MVI), neural invasion identification (NI), pan-cancer identification across 32 types (PanCancer), organ classification (OC), tissue classification (TC), tumor-infiltrating lymphocyte identification (TIL), microsatellite instability detection in colorectal cancer (MSI), and seven-class gastric lesion recognition (GLR-7). The accuracy of each model on these tasks is detailed in Tab. 1.
Slide-level experiments included not only the same subtyping and grading tasks as the patch-level but also additional tasks, such as lymph node metastasis diagnosis (LNM), HCC prognosis prediction (HCC-P), and colorectal cancer prognosis prediction (CRC-P), using WSI thumbnails as image input. The accuracy of each model on slide-level tasks is presented in Tab. 2.
It can be observed that OmniPath achieves the best performance across all patch-level and slide-level pathological diagnosis tasks. For most cancer subtype classification and grading tasks, OmniPath achieves accuracy rates exceeding 70%, and in many cases, surpassing 90%. In contrast, the accuracy rates of other models are generally below 70%. This demonstrates that OmniPath is more suitable for clinical applications to assist pathologists in diagnosis. Furthermore, OmniPath also exhibits strong recognition capabilities for features such as microvascular invasion, neural invasion, and tumor-infiltrating lymphocytes. This significantly reduces the extensive effort required by pathologists to meticulously examine detailed pathological lesions during slide review. Moreover, OmniPath outperforms Quilt-LLaVA (FT), demonstrating the effectiveness of the MGFE and PGFC strategies.
Table 1: Accuracy (%) comparison on patch-level pathology diagnostic tasks.
@@ 5.2 Comparison on Pathology Diagnostic Tasks
Table 2: Accuracy (%) comparison on slide-level pathology diagnostic tasks.
@@ 5.3 Comparison on Zero-Shot Classification Tasks
To evaluate the clinical generalization capability of OmniPath, we employed a zero-shot classification paradigm, testing on several widely-used academic pathology datasets that were not included in the training set. The evaluation covered two levels: patch-level tasks using the CCRCC [CCRCC], MHIST [MHIST], and NCT-CRC [NCT-CRC] datasets, and slide-level tasks using the PANDA [PANDA], DHMC [DHMC], and CAMELYON17 [Camelyon17] datasets. Using a closed-ended question-answering approach, the model was required to classify images into predefined categories specific to each dataset. The performance comparison of all models on these zero-shot test sets is presented in Tab. 4.
It is shown that OmniPath consistently outperforms other models across both patch-level and slide-level datasets in zero-shot classification tasks, highlighting its strong generalization ability in pathological image analysis. Notably, on the PANDA and CAMELYON17 slide-level datasets, OmniPath achieved the highest accuracy rates of 79.15% and 59.33%, respectively, which significantly surpasses the performance of other models. This superior performance in zero-shot classification indicates OmniPath’s robustness in handling diverse pathological image data and reinforces its potential for clinical applications where labeled training data may be limited.
Table 3: Accuracy (%) on zero-shot classification.
@@
Table 4: Performance comparison on referring segmentation tasks. Dice coefficient (%) is used as the evaluation metric.
Method Slide-Level Cancer Region Tissue Structure
@@
Quilt-LLaVA (FT) 62.10 11.95 23.82 4.82 18.67 16.77 24.32 3.25 4.01
OmniPath 95.51 93.60 85.11 50.01 93.03 74.64 89.23 62.14 56.94
Table 5: Performance comparison on pathology referring detection tasks.
Method Metric Slide-Level Cancer Region Tissue Structure Cell Nucleus
HCC ICC RCC GBM LUAD BC LD VD ND MoNuSeg NuCLS PanNuke
Quilt-LLaVA (FT) F1-Score (%) 84.64 77.27 63.33 14.38 63.33 11.76 32.76 69.81 15.03 14.27 5.63 10.17
IoU (%) 78.14 78.62 73.38 63.69 69.86 87.26 67.03 69.03 66.62 63.05 19.59 26.73
F1-Score (%) 92.13 100.00 90.91 40.85 90.91 41.38 82.70 89.52 72.50 83.98 34.30 45.30
OmniPath IoU (%) 91.97 91.85 89.59 73.85 87.96 72.75 87.34 83.00 79.75 79.65 44.98 59.99
@@ 5.4 Detection and Segmentation Performance
In pathological diagnosis, detection and segmentation tasks are as critical as classification tasks. Since existing pathology LVLMs lack detection and segmentation capabilities, we only compare with Quilt-LLaVA (FT). To enable Quilt-LLaVA to perform detection and segmentation, we used the same training dataset as OmniPath and converted the segmentation masks in the dataset into corresponding polygons with up to 50 vertices each. We then had Quilt-LLaVA generate the vertex coordinate sequences of the polygons as text sequences. The detection tasks cover various cancer region identifications, including HCC, ICC, RCC, glioblastoma (GBM), lung adenocarcinoma (LUAD), and bladder cancer (BC). Additionally, it involves detecting tissue structures such as lymph nodes (LD), vessels (VD), and nerves (ND), as well as cell nuclei detection in datasets like MoNuSeg [MoNuSeg] (without categories), NuCLS [NuCLS], and PanNuke [PanNuke] (with categories). The segmentation tasks not only include the segmentation of cancer regions covered by the detection tasks but also nerve segmentation (NS), nerve invasion region segmentation (NIS), and lymph node metastasis segmentation (LNMS). Detection tasks are evaluated using F1-score and IoU, while segmentation tasks are assessed using the Dice coefficient. Tab. 4 and Tab. 5 show the segmentation and detection performance, respectively.
The “Cancer Region” detection and segmentation tasks are slide-level, and all remaining tasks are patch-level. OmniPath outperforms Quilt-LLaVA (FT) in both detection and segmentation. While OmniPath’s performance may not surpass specialized smaller models on certain tasks (see Appendix C), it offers a unique advantage in inferential capabilities. As shown in Fig. 1, OmniPath can accurately detect cancer cell nuclei within blood vessels according to instructions (see Appendix E for quantitative results)—an ability that current specialized small models find challenging to achieve.